# A Unified Debiasing Approach for Vision-Language Models across Modalities and Tasks

### v1

@@ 1 Introduction
Vision-Language Models (VLMs) have revolutionized the way we handle multimodal tasks by enabling simultaneous processing of text and image data. Models such as CLIP Radford et al. (2021) and XVLM Zeng et al. (2021) serve as foundation models Xu et al. (2024) for various downstream tasks demonstrating the remarkable versatility of these systems such as zero-shot classification and text-to-image retrieval. Additionally, models like BLIP Li et al. (2022) and CoDi Tang et al. (2024) enhance this spectrum by facilitating tasks such as image captioning and text-to-image generation. These examples highlight the diverse capabilities of VLMs in adapting to various specific multimodal interactions.
Despite their remarkable capabilities, VLMs have a critical bias issues, often skewing the model outputs in ways that reflect societal stereotypes Janghorbani and De Melo (2023) such as gender Zhou et al. (2022) or racial Zhao et al. (2021) biases in assigning professions or describing scenarios. For example, studies have identified biases in multi-class zero-shot classification Gustafson et al. (2023), where models might disproportionately associate certain professions with specific genders. Similarly, biases in text-to-image retrieval Hamidieh et al. (2023); Wang et al. (2022) can lead to the preferential retrieval of images that reinforce stereotypical narratives. The implications of these biases extend to image captioning Zhao et al. (2021); Hirota et al. (2023) and text-to-image generation Cho et al. (2023), where the descriptive and generative capacities of VLMs may perpetuate and even amplify existing societal prejudices. These issues highlight a critical need for effective debiasing strategies that can ensure the fairness of VLMs applications.
@@
Given the bias issues in the realm of VLMs, debiasing approaches have been proposed. However, the existing debiasing frameworks for VLMs tend to focus on specific modalities and tasks. For example, Hirota et al. (2023) and Kim et al. (2023a) are tailored to particular tasks such as image captioning or text-to-image generation, respectively. On the other hand, while some debiasing strategies are originally designed for mitigating bias in the image or text encoder’s output for retrieval task, like DeAR Seth et al. (2023) and CLIP-clip Wang et al. (2021), they offer potential for broader application across various modalities and tasks. For instance, DeAR can extend the use of adversarial training to neutralize biases in the decoder’s intermediate representations by aligning them near the decision boundary for sensitive attribute by deceiving the attribute classifier. Yet, this method’s effectiveness is hampered by the intrinsic complexities and sensitivities of adversarial training techniques, which are highly dependent on precise hyperparameter tuning. Similarly, CLIP-clip could potentially adapt to various tasks by pruning features in the frozen representation that exhibit high mutual information with sensitive attributes. However, this approach does not consider dependencies between feature columns. Furthermore, the process of zeroing out pruned features might inadvertently distort the semantic meaning of embeddings, leading to a loss in the quality and relevance of the output. Moreover, Prompt-Debias Chuang et al. (2023), which relies on pre-defined text prompts to debias the text encoder, is limited to text encoders and cannot be applied to text decoders or other components.
@@
To overcome these limitations, we propose Selective Feature Imputation for Debiasing (SFID) incorporating feature pruning and low confidence imputation (LCI). Unlike existing methods, this approach can be seamlessly integrated into various parts of VLMs, whether applied to encoders, decoders, or both, enhancing its utility in diverse contexts, while effecitvely maintain the dimensionalty and semantical meaning of debiased representation. In details, by utilizing feature selection techniques such as RandomForest Breiman (2001) to identify gender-specific (or race) biases within the frozen representation, SFID prunes and subsequently replace bias-causing features with bias-free representation obtained from ambiguous samples identified by RandomForest, thereby maintaining the semantic content while effectively reducing bias.
@@
Furthermore, SFID eliminates the need for costly retraining of pre-trained VLMs and does not require paired text-image datasets. It simply utilizes datasets with sensitive attributes in individual images or texts for debiasing. For instance, to debias gender bias in VLMs, datasets like FairFace Kärkkäinen and Joo (2019) for image inputs and Bias-in-Bios De-Arteaga et al. (2019) for text inputs are employed. This approach not only maintains the utility and efficiency of VLMs but also broadens their applicability across varied scenarios, thereby enhancing fairness without compromising performance.
@@ 2 Related Work
2.1 Bias evluation in VLMs
Recent research has focused significantly on identifying and evaluating bias in VLMs. Agarwal et al. (2021) and Wolfe et al. (2022) observed that CLIP (Radford et al., 2021) embeddings exhibit substantial racial and gender biases. Additionally, Chuang et al. (2023) highlighted that text prompts can capture spurious features in visual representations, exacerbating the bias. Kim et al. (2023b) further investigated how the association between sensitive attributes and specific keywords contributes to bias issues in downstream tasks. These biases lead to unfair outcomes in zero-shot binary classification and text-to-image retrieval, as noted by (Dehdashtian et al., 2024; Hamidieh et al., 2023). Moreover, Slyman et al. (2024) extend the bias in zero-shot classification in multi-class setting using the FACET dataset Gustafson et al. (2023) and its evaluation metric. Seth et al. (2023) suggested a new metric for fairness in text-to-image retrieval, considering the gender distribution in the query set. Beyond zero-shot classification and text-to-image retrieval, other downstream tasks also reveal biases. For instance, Hirota et al. (2023) and Zhao et al. (2021) investigated bias in image captioning, where specific genders or races are disproportionately represented leading to generate biased caption. Similarly, Cho et al. (2023) raised concerns about bias in text-to-image generation and suggested evaluation methods. Finally, Sathe et al. (2024) emphasized the need for a unified evaluation approach for various downstream tasks to address these pervasive biases comprehensively.
@@ 2.2 Debiasing VLMs
As bias issues have arisen, many debiasing methods have been proposed. Zhang and Ré (2022) trained an adapter on frozen representations to debias spurious correlations in zero-shot binary classification. Additionally, Chuang et al. (2023); Adila et al. (2023), and Berg et al. (2022) suggested methods to manipulate input prompts or text tokens for debiasing spurious correlations in VLMs’ encoders, covering downstream tasks such as zero-shot binary classification and text-to-image retrieval. For image captioning, Hirota et al. (2023) proposed a fine-tuning method for both the encoder and decoder to mitigate biases. For text-to-image generation, Kim et al. (2023a) recommended a de-stereotyping prompt design to address biases. Some methods, such as DeAR Seth et al. (2023) and CLIP-clip Wang et al. (2021), manipulate frozen representations without training the entire model and can be generalized across various tasks, as discussed in Section 4. Dehdashtian et al. (2024) also aimed to debias frozen representations from both the image and text encoders, though this approach requires class labels and text-image pair datasets, which are challenging to define and obtain across various downstream tasks. To address these limitations, we propose the first unified debiasing method for VLMs across various downstream tasks, which demonstrates outstanding performance compared to the extensions of other debiasing methods, such as DeAR and CLIP-clip.
@@ 3.1 Zero-shot Classification
Multi-class zero-shot classification leverages the capability of VLMs that train image and text encoders jointly. Specifically, the predicted class is determined by providing text prompts about classes to the text encoder and selecting the class with the highest cosine similarity to the encoded image. For instance, the text prompt “a photo of a/an [CLASS NAME]” is used for the zero-shot classification. Bias issues arise from the difference in accuracy between genders for a given class. For example, in the class Carpenter, the accuracy for male carpenter images and female carpenter images might differ, as VLMs are likely to associate the concept of “male” more strongly with “carpenter.” Therefore, bias is defined by the average demographic disparity, which is determined by the gender disparity in recall for each class following Gustafson et al. (2023),
@@
where $\hat{Y}$ is the predicted class, and $k\in\mathcal{K}$ represents each class in the multi-class classification, and $a\in\{0,1\}$ is the sensitive attribute. The overall accuracy is used as an evaluation metric for the performance of VLMs. Lower $\Delta DP$ indicates fair classification across the sensitive attribute, while higher overall accuracy denotes higher classification ability. As a baseline for zero-shot classification, CLIP Radford et al. (2021) with ResNet-50 He et al. (2016) and ViT-B/32 Dosovitskiy et al. (2020), and XVLM Zeng et al. (2021) are adopted. We utilize the FACET Gustafson et al. (2023) dataset, which includes 49,551 images across 52 classes with gender sensitive attribute.
@@ 3.2 Text-to-Image Retrieval
The evaluation metric for fairness in text-to-image retrieval is defined as $Skew@M$, as suggested in Seth et al. (2023). Let $p_{a}=N_{a}/N$ for $a\in\{0,1\}$, where $N_{a}$ is the number of images for each gender in the original dataset and $N$ is the total number of images. For $M$ retrieved images, calculate $\hat{p}_{a}=M_{a}/M$, where $M_{a}$ is the number of images for each sensitive attribute in the retrieved set. The metric is then defined as $Skew_{a}=\log(\hat{p}_{a}/p_{a})$ for each sensitive attribute $a$ indicating if a particular gender is retrieved more frequently. The final evaluation metric, $Skew$, is defined by averaging the maximum Skew values over the set of text prompts $\mathcal{T}$,
@@
The performance metric for text-to-image retrieval is defined as $Recall@K$, which measures the probability of the ground truth image being among the top $K$, where $K\in\{1,5,10\}$. Thus, lower $Skew$ and higher $Recall$ are desired. We use the same baselines: CLIP ResNet-50, CLIP ViT-B/32, and XVLM. In the experiments, we utilize the Flickr30K Young et al. (2014) dataset which includes ground truth captions and gender attributes. After modifying the ground truth captions to be gender-neutral, we select 1,000 images for testing Wang et al. (2021) and retrieve the top 100 images for each caption for bias measure.
@@ 3.3 Image Captioning
Image captioning aims to generate a caption given an image. Fairness issues arise from differences between the gender mentioned in the caption and the ground truth gender of the subjects in the image Hirota et al. (2023), as VLMs may prefer certain genders for particular subjects. For example, as shown in Figure 1, image captioning models tend to associate contexts and genders, such as (athlete, male) and (long hair, female).
@@
where the misclassification rate for each gender is defined as $MR_{M}=\frac{1}{|\mathcal{M}|}\sum_{k\in\mathcal{M}}I_{k}$, $MR_{F}=\frac{1}{|\mathcal{F}|}\sum_{k\in\mathcal{F}}I_{k}$, and $MR_{O}=\frac{1}{|\mathcal{D}|}\sum_{k\in\mathcal{D}}I_{k}$, with $M$, $F$, and $O$ indicating male, female, and overall, respectively. Although the overall misclassification rate is used in Hirota et al. (2023), it cannot perfectly reflect fairness. For example, in two different situations where $(MR_{M},MR_{F},MR_{O})$ are (3.0%, 3.0%, 3.0%) and (0.0%, 6.0%, 3.0%), respectively, the overall misclassification rates are the same, but the rates for each gender are not fair. To address this, we derive a Composite Misclassification Rate, defined as $MR_{C}=\sqrt{MR_{O}^{2}+(MR_{F}-MR_{M})^{2}}$, which can be minimized when both the overall misclassification rate and the disparity in misclassification rates between genders are low.
On the other hand, the caption’s quality is measure by METEOR Banerjee and Lavie (2005) and SPICE Anderson et al. (2016). METEOR measures the balance between precision and recall of n-grams in generated captions, incorporating synonyms, while SPICE focuses on the semantic content of captions by comparing sets of propositional semantic tuples extracted from candidate and reference captions. (See Appendix B for details.) Considering fair image captioning, the evaluation metric for caption quality should account for both the ground truth caption and a neutral caption.
@@
In image captioning, a lower $MR_{C}$ score and higher METEOR and SPICE scores are preferable. In the experiment, due to ambiguity in images containing more than one person, such images are excluded from the evaluation. We adopt two image captioning models, CLIP-cap Mokady et al. (2021) and BLIP Li et al. (2022). The MS-COCO dataset Chen et al. (2015) is used as the query dataset, which consists of 10,780 images with 5 ground truth captions for each image.
@@ 3.4 Text-to-Image Generation
VLMs can be used for text-to-image generation, creating an image given a text prompt. However, bias issues in text-to-image generation arise in two ways: generating an image with the wrong gender despite of a specific gender is given in the prompt, or generating a biased gendered image for a neutral text prompt. We utilize the input prompt from Cho et al. (2023) to generate images for captions: “a photo of $\mathcal{G}$ who works as a/an $\mathcal{P}$,” where gender set $\mathcal{G}=$a man, a woman, a person} and profession set $\mathcal{P}=$accountant, engineer, $\cdots$}. We use 83 professions as in Cho et al. (2023).
For the evaluation, we first detect the gender of the generated image by utilizing BLIP-2 Li et al. (2023) by asking “Does the person look like a male or a female?”. For gender-specific prompts, a fair text-to-image generation model should produce exact gender images irrespective of the profession. Thus, we measure the mismatch rate: $\frac{1}{|\mathcal{D}|}\sum_{k\in\mathcal{D}}I_{k}$, where $\mathcal{D}$ is a set of prompts for either the male or female subgroup. We adapt the composite misclassification rate, $MR_{C}$, as the same as the image captioning to verify a balanced outcome between mismatch rates for each gender. For neutral prompts, a fair model should produce an equal number of images of each gender across all professions. The bias metric thus measures the skewed distribution of the model:
@@
where $N_{p,m}$ and $N_{p,f}$ are the numbers of detected genders for each profession, and $C=10$ is the number of generation for each prompt. In text-to-image generation, lower values of both $MR_{C}$ and $Skew$ indicate fairness. We use SDXL Podell et al. (2023) and CoDi Tang et al. (2024) as baselines for text-to-image generation.

### v2

@@ 1 Introduction
Vision-Language Models (VLMs) have revolutionized the way we handle multimodal tasks by enabling simultaneous processing of text and image data. Models such as CLIP [31] and XVLM [43] serve as foundation models [41] for various downstream tasks demonstrating the remarkable versatility of these systems such as zero-shot classification and text-to-image retrieval. Additionally, models like BLIP [25] and CoDi [37] enhance this spectrum by facilitating tasks such as image captioning and text-to-image generation. These examples highlight the diverse capabilities of VLMs in adapting to various specific multimodal interactions.
Despite their remarkable capabilities, VLMs have a critical bias issues, often skewing the model outputs in ways that reflect societal stereotypes [21] such as gender [46] or racial [45] biases in assigning professions or describing scenarios. For example, studies have identified biases in multi-class zero-shot classification [17], where models might disproportionately associate certain professions with specific genders. Similarly, biases in text-to-image retrieval [18, 39] can lead to the preferential retrieval of images that reinforce stereotypical narratives. The implications of these biases extend to image captioning [45, 20] and text-to-image generation [10], where the descriptive and generative capacities of VLMs may perpetuate and even amplify existing societal prejudices. These issues highlight a critical need for effective debiasing strategies that can ensure the fairness of VLMs applications.
@@
Given the bias issues in the realm of VLMs, debiasing approaches have been proposed. However, the existing debiasing frameworks for VLMs tend to focus on specific modalities and tasks. For example, [20] and [23] are tailored to particular tasks such as image captioning or text-to-image generation, respectively. On the other hand, while some debiasing strategies are originally designed for mitigating bias in the image or text encoder’s output for retrieval task, like DeAR [34] and CLIP-clip [38], they offer potential for broader application across various modalities and tasks. For instance, DeAR can extend the use of adversarial training to neutralize biases in the decoder’s intermediate representations by aligning them near the decision boundary for sensitive attribute by deceiving the attribute classifier. Yet, this method’s effectiveness is hampered by the intrinsic complexities and sensitivities of adversarial training techniques, which are highly dependent on precise hyperparameter tuning. Similarly, CLIP-clip could potentially adapt to various tasks by pruning features in the frozen representation that exhibit high mutual information with sensitive attributes. However, this approach does not consider dependencies between feature columns. Furthermore, the process of zeroing out pruned features might inadvertently distort the semantic meaning of embeddings, leading to a loss in the quality and relevance of the output. Moreover, Prompt-Debias [11], which relies on pre-defined text prompts to debias the text encoder, is limited to text encoders and cannot be applied to text decoders or other components.
@@
To overcome these limitations, we propose Selective Feature Imputation for Debiasing (SFID) incorporating feature pruning and low confidence imputation (LCI). Unlike existing methods, this approach can be seamlessly integrated into various parts of VLMs, whether applied to encoders, decoders, or both, enhancing its utility in diverse contexts, while effecitvely maintain the dimensionalty and semantical meaning of debiased representation. In details, by utilizing feature selection techniques such as RandomForest [8] to identify gender-specific (or race) biases within the frozen representation, SFID prunes and subsequently replace bias-causing features with bias-free representation obtained from ambiguous samples identified by RandomForest, thereby maintaining the semantic content while effectively reducing bias.
@@
Furthermore, SFID eliminates the need for costly retraining of pre-trained VLMs and does not require paired text-image datasets. It simply utilizes datasets with sensitive attributes in individual images or texts for debiasing. For instance, to debias gender bias in VLMs, datasets like FairFace [22] for image inputs and Bias-in-Bios [12] for text inputs are employed. This approach not only maintains the utility and efficiency of VLMs but also broadens their applicability across varied scenarios, thereby enhancing fairness without compromising performance.
@@ 2 Related Work
2.1 Bias Evaluation in VLMs
Recent research has focused significantly on identifying and evaluating bias in VLMs. Agarwal et al. [2] and Wolfe et al. [40] observed that CLIP [31] embeddings exhibit substantial racial and gender biases. Additionally, Chuang et al. [11] highlighted that text prompts can capture spurious features in visual representations, exacerbating the bias. Kim et al. [24] further investigated how the association between sensitive attributes and specific keywords contributes to bias issues in downstream tasks. These biases lead to unfair outcomes in zero-shot binary classification and text-to-image retrieval, as noted by [13, 18]. Moreover, Slyman et al. [35] extend the bias in zero-shot classification in multi-class setting using the FACET dataset [17] and its evaluation metric. Seth et al. [34] suggested a new metric for fairness in text-to-image retrieval, considering the gender distribution in the query set. Beyond zero-shot classification and text-to-image retrieval, other downstream tasks also reveal biases. For instance, Hirota et al. [20] and Zhao et al. [45] investigated bias in image captioning, where specific genders or races are disproportionately represented leading to generate biased caption. Similarly, Cho et al. [10] raised concerns about bias in text-to-image generation and suggested evaluation methods. Finally, Sathe et al. [32] emphasized the need for a unified evaluation approach for various downstream tasks to address these pervasive biases comprehensively.
@@ 2.2 Debiasing VLMs
As bias issues have arisen, many debiasing methods have been proposed. Zhang and Ré [44] trained an adapter on frozen representations to debias spurious correlations in zero-shot binary classification. Additionally, Chuang et al. [11], Adila et al. [1], and Berg et al. [6] suggested methods to manipulate input prompts or text tokens for debiasing spurious correlations in VLMs’ encoders, covering downstream tasks such as zero-shot binary classification and text-to-image retrieval. For image captioning, Hirota et al. [20] proposed a fine-tuning method for both the encoder and decoder to mitigate biases. For text-to-image generation, Kim et al. [23] recommended a de-stereotyping prompt design to address biases. Some methods, such as DeAR [34] and CLIP-clip [38], manipulate frozen representations without training the entire model and can be generalized across various tasks, as discussed in Section 4. Dehdashtian et al. [13] also aimed to debias frozen representations from both the image and text encoders, though this approach requires class labels and text-image pair datasets, which are challenging to define and obtain across various downstream tasks. To address these limitations, we propose the first unified debiasing method for VLMs across various downstream tasks, which demonstrates outstanding performance compared to the extensions of other debiasing methods, such as DeAR and CLIP-clip.
@@ 3.1 Zero-shot Classification
Multi-class zero-shot classification leverages the capability of VLMs that train image and text encoders jointly. Specifically, the predicted class is determined by providing text prompts about classes to the text encoder and selecting the class with the highest cosine similarity to the encoded image. For instance, the text prompt “a photo of a/an [CLASS NAME]” is used for the zero-shot classification. Bias issues arise from the difference in accuracy between genders for a given class. For example, in the class Carpenter, the accuracy for male carpenter images and female carpenter images might differ, as VLMs are likely to associate the concept of “male” more strongly with “carpenter.” Therefore, bias is defined by the average demographic disparity, which is determined by the gender disparity in recall for each class following [17],
@@
where $\hat{Y}$ is the predicted class, and $k\in\mathcal{K}$ represents each class in the multi-class classification, and $a\in\{0,1\}$ is the sensitive attribute. The overall accuracy is used as an evaluation metric for the performance of VLMs. Lower $\Delta DP$ indicates fair classification across the sensitive attribute, while higher overall accuracy denotes higher classification ability. As a baseline for zero-shot classification, CLIP [31] with ResNet-50 [19] and ViT-B/32 [15], and XVLM [43] are adopted. We utilize the FACET [17] dataset, which includes 49,551 images across 52 classes with gender sensitive attribute.
@@ 3.2 Text-to-Image Retrieval
The evaluation metric for fairness in text-to-image retrieval is defined as $Skew@M$, as suggested in [34]. Let $p_{a}=N_{a}/N$ for $a\in\{0,1\}$, where $N_{a}$ is the number of images for each gender in the original dataset and $N$ is the total number of images. For $M$ retrieved images, calculate $\hat{p}_{a}=M_{a}/M$, where $M_{a}$ is the number of images for each sensitive attribute in the retrieved set. The metric is then defined as $Skew_{a}=\log(\hat{p}_{a}/p_{a})$ for each sensitive attribute $a$ indicating if a particular gender is retrieved more frequently. The final evaluation metric, $Skew$, is defined by averaging the maximum Skew values over the set of text prompts $\mathcal{T}$,
@@
The performance metric for text-to-image retrieval is defined as $Recall@K$, which measures the probability of the ground truth image being among the top $K$, where $K\in\{1,5,10\}$. Thus, lower $Skew$ and higher $Recall$ are desired. We use the same baselines: CLIP ResNet-50, CLIP ViT-B/32, and XVLM. In the experiments, we utilize the Flickr30K [42] dataset which includes ground truth captions and gender attributes. After modifying the ground truth captions to be gender-neutral, we select 1,000 images for testing [38] and retrieve the top 100 images for each caption for bias measure.
@@ 3.3 Image Captioning
Image captioning aims to generate a caption given an image. Fairness issues arise from differences between the gender mentioned in the caption and the ground truth gender of the subjects in the image [20], as VLMs may prefer certain genders for particular subjects. For example, as shown in Figure 1, image captioning models tend to associate contexts and genders, such as (athlete, male) and (long hair, female).
@@
where the misclassification rate for each gender is defined as $MR_{M}=\frac{1}{|\mathcal{M}|}\sum_{k\in\mathcal{M}}I_{k}$, $MR_{F}=\frac{1}{|\mathcal{F}|}\sum_{k\in\mathcal{F}}I_{k}$, and $MR_{O}=\frac{1}{|\mathcal{D}|}\sum_{k\in\mathcal{D}}I_{k}$, with $M$, $F$, and $O$ indicating male, female, and overall, respectively. Although the overall misclassification rate is used in [20], it cannot perfectly reflect fairness. For example, in two different situations where $(MR_{M},MR_{F},MR_{O})$ are (3.0%, 3.0%, 3.0%) and (0.0%, 6.0%, 3.0%), respectively, the overall misclassification rates are the same, but the rates for each gender are not fair. To address this, we derive a Composite Misclassification Rate, defined as $MR_{C}=\sqrt{MR_{O}^{2}+(MR_{F}-MR_{M})^{2}}$, which can be minimized when both the overall misclassification rate and the disparity in misclassification rates between genders are low.
On the other hand, the caption’s quality is measure by METEOR [4] and SPICE [3]. METEOR measures the balance between precision and recall of n-grams in generated captions, incorporating synonyms, while SPICE focuses on the semantic content of captions by comparing sets of propositional semantic tuples extracted from candidate and reference captions. (See Appendix B for details.) Considering fair image captioning, the evaluation metric for caption quality should account for both the ground truth caption and a neutral caption.
@@
In image captioning, a lower $MR_{C}$ score and higher METEOR and SPICE scores are preferable. In the experiment, due to ambiguity in images containing more than one person, such images are excluded from the evaluation. We adopt two image captioning models, CLIP-cap [27] and BLIP [25]. The MS-COCO dataset [9] is used as the query dataset, which consists of 10,780 images with 5 ground truth captions for each image.
@@ 3.4 Text-to-Image Generation
VLMs can be used for text-to-image generation, creating an image given a text prompt. However, bias issues in text-to-image generation arise in two ways: generating an image with the wrong gender despite of a specific gender is given in the prompt, or generating a biased gendered image for a neutral text prompt. We utilize the input prompt from [10] to generate images for captions: “a photo of $\mathcal{G}$ who works as a/an $\mathcal{P}$,” where gender set $\mathcal{G}=$a man, a woman, a person} and profession set $\mathcal{P}=$accountant, engineer, $\cdots$}. We use 83 professions as in [10].
For the evaluation, we first detect the gender of the generated image by utilizing BLIP-2 [26] by asking “Does the person look like a male or a female?”. For gender-specific prompts, a fair text-to-image generation model should produce exact gender images irrespective of the profession. Thus, we measure the mismatch rate: $\frac{1}{|\mathcal{D}|}\sum_{k\in\mathcal{D}}I_{k}$, where $\mathcal{D}$ is a set of prompts for either the male or female subgroup. We adapt the composite misclassification rate, $MR_{C}$, as the same as the image captioning to verify a balanced outcome between mismatch rates for each gender. For neutral prompts, a fair model should produce an equal number of images of each gender across all professions. The bias metric thus measures the skewed distribution of the model:
@@
where $N_{p,m}$ and $N_{p,f}$ are the numbers of detected genders for each profession, and $C=10$ is the number of generation for each prompt. In text-to-image generation, lower values of both $MR_{C}$ and $Skew$ indicate fairness. We use SDXL [28] and CoDi [37] as baselines for text-to-image generation.